# The Need for Benchmarks to Advance AI-Enabled Player Risk Detection in Gambling

### v1

@@ 1 Introduction
Gambling is a significant and growing form of entertainment, participated in by millions globally. While most consumers enjoy gambling as a leisure activity, a small percentage of participants experience serious harms, while a larger proportion may experience some subclinical negative effects (Tran2024). The industry’s rapid expansion, driven largely by digitization and the global availability of online and mobile products, has led to a corresponding increase in efforts to mitigate these potential harms (Ghaharian2022). These collective efforts are commonly referred to as Responsible Gambling (RG) [^1] We use the term Responsible Gambling (RG) throughout this paper as it is commonly used to describe these practices. We acknowledge the ongoing and important discourse contrasting the RG framework to safer gambling or public health approaches, but a full discussion of that debate is beyond the scope of this paper..
@@
RG strategies have evolved as a direct result of gambling’s digitization, and include targeted RG messaging, immediate access to gambling help resources, and automated risk detection systems (Shaffer2020). These technological player protection measures can be divided roughly into two types: user-initiated measures and automated operator-initiated strategies. The former includes tools that users can employ, usually voluntarily, to control their gambling: self-limiting features, reality checks, deposit limits, cool-off periods, and self-exclusion. However, few users take advantage of these features (Ladouceur2016; Nelson2008), and there is debate about whether those who do are the players who need them most (catania2021understanding).
@@
While automated, operator-initiated strategies can also take many forms, including universal approaches like deposit limits, there has been a movement in the past decade toward more tailored interventions (heirene2025setting). Universal strategies risk either disrupting the activity of those who do not need protections or, if implemented in a way that does not disrupt most users’ activity, being set so conservatively that many of those who need them do not trigger them (broda2008optimal). On the other hand, targeted strategies can, in theory, use player information to better determine when and to whom to deploy protections.
@@
As such, data-driven RG models have become the holy grail of automated operator-initiated RG solutions (Ghaharian2022). These systems often leverage artificial intelligence (AI) – more specifically, machine learning (ML) – to analyze players’ gambling behavior and detect patterns associated with harm to trigger interventions. This approach has attracted substantial interest from both academia and industry stakeholders, with several regulators mandating the use of data-driven systems for player risk detection, operators developing proprietary models, and specialist technology vendors emerging to offer commercial solutions.
@@
The paper is presented as follows: We first provide a brief background on the use of ML for player risk detection, followed by a discussion of the current challenges in evaluating these systems. We then introduce the concept of Performance Benchmarking and briefly review its application in other domains. Finally, we present a conceptual, discussion-starting framework for a player risk detection benchmarking suite, outlining its core components and the requirements for its successful implementation. Our hope is to stimulate discussion and thought for what such a framework would require and what it could achieve.
@@ 1.1 Machine Learning for Player Risk Detection
A ML approach to player risk detection typically uses information about gambling patterns and player characteristics from gambling records (behavioral markers) to forecast potential gambling-related problems. This is similar to the use of biomarkers in medicine, where objectively measured characteristics serve as indicators of disease processes (biomarkers2001biomarkers). Medicine also uses behavioral markers, for example, in early detection of Alzheimer’s disease (gomar2011utility).
@@
ML models can target players according to markers defined by both their gambling patterns and changes in those patterns, using this information to designate differing levels of potential risk. Most importantly, these models can provide a guide for matching responses to individuals, optimizing the opportunity for favorable outcomes. Very basic best practices for developing such models include establishing that the markers differ in the groups to be distinguished (e.g., individuals who exhibit problematic gambling behavior and those who gamble without problems), developing a predictive model that integrates markers to predict the outcome in question, and then establishing the model’s performance in samples separate from those used to develop the model (Morrow2008; vanderstichele2008).
@@
One distinct feature of ML is its ability to learn complex, nonlinear relationships in data, making it highly suited for identifying at-risk play given the individual variability in gambling behavior (hastie2009elements). However, complex ML solutions encounter some of the same challenges as the most basic models, such as establishing accurate outcome measures and training data and developing models that can generalize and replicate beyond the idiosyncrasies of a specific dataset (i.e., distinguishing the signal from the noise: silver2012signal; also see Shaffer2020).
@@
Early player risk models re-purposed data from loyalty smart cards used in physical gambling venues (e.g., casinos) (Schellinck2011). The shift to online gambling, however, significantly simplified behavioral tracking. Unlike physical venues that depended on a player inserting a card, online platforms and mobile apps inherently capture comprehensive player data. This seamless data collection fueled a new research area, leading to at least 40 academic studies and several literature reviews (we refer the interested reader to: Ghaharian2022; Delfabbro2023; MARIONNEAU2025100644). Alongside this academic work, commercial models were developed. A recent review by MARIONNEAU2025100644 identified 11 commercial models, including six in-house systems from operators (e.g., Kindred, Veikkaus) and five from third-party vendors (e.g., Mindway, Openbet).
@@
These models vary widely in their methods, employing a variety of unsupervised and supervised algorithms, and in the number and type of input markers they use. As MARIONNEAU2025100644 note, these markers are often unclear or inconsistently reported in the literature, and the number of markers used in a given model range from fewer than 10 to over 100. ghaharian2025ai recently mapped the markers in the academic literature, identifying 65 “low-level” categories (such as bet amount and net loss) which can be grouped into five “high-level” categories: play (betting behavior), payment (financial transactions, i.e., deposits and withdrawals), engagement (when, how often, and how broadly players engage with gambling), RG tool use, and profile information (e.g., registration date, demographic information).
@@ 2 Current Problems
It is appropriate to hold ML to high standards, given that policy increasingly references it in supporting harm reduction (bedford2024high). Moreover, such standards align with the sensible and realistic performance beliefs that gamblers themselves may hold (murch2025preferences). Several challenges limit the potential for ML in this context.
@@ 2.1 Broad Challenges for Stakeholders
ML, as with all identification and intervention design, involves trade-offs between sensitivity and precision, i.e., where to place the balance between failing to identify gamblers at risk and falsely identifying many gamblers with negligible risk levels (Murch2023). These trade-offs can be improved with more sophisticated models but cannot be fully resolved in real-world settings. In the absence of precision, an ultra-low specificity model simply provides RG interventions to all players, which may undermine trust in RG programs and potentially lead to substitution by players towards other operators or products.
@@
As models grow more complex, they typically become less transparent with respect to how any individual decision has been reached (sarkar2016accuracy). This is often referred to as the “black box” problem, where the model’s internal logic is so intricate that it becomes opaque even to its creators, making it difficult to audit or explain a specific outcome (vonEschenbach2021). Explainability, or the ability to identify how a model arrives at its prediction, is important to facilitate users’ trust in a system and ensure the model is operating as expected (sarkar2016accuracy). In addition, explainability has practical relevance: it enables operators to activate appropriate interventions and allows players to reflect on behavioral dimensions that have been identified as increasing their risk.
@@
Another issue is that models can only identify patterns in the data they possess. Consequently, identical gambling behaviors can represent very different risk-levels depending on variables not included in a data set, such as an individual player’s life circumstances or financial context. For example, the potential impact of a specific monetary loss is difficult to contextualize without information about an individual’s financial circumstances (e.g., income, debt exposure, etc.). Moreover, harms are not purely financial. A significant amount of time spent gambling could be a harmless hobby for one person, but for someone suffering psychological distress, it could be a coping mechanism leading to other harms like relationship disruption or missing work. Much of this relevant contextual information simply cannot be collected by operators under current social norms. This missing context is why seemingly identical profiles based on available data could have different risks and, as the gamblingcommission2023risk has noted, why high-risk players exist at all stake levels.
@@
There are also challenges related to AI ethics generally, such as risks of bias in datasets or models (percy2020lessons), data collection protocols, and ensuring human oversight for high-consequence decisions. While a full discussion of AI ethics in gambling is beyond this paper’s scope (discussions available in binesh2025identifying; ghaharian2025aiethics; Percy2022), we note the trend towards translating ethical principles into regulation (e.g., the EU AI Act approved in May 2024).
@@ 2.2 Specific Challenges for Procuring or Commissioning Models
Given the quantitative nature of data science, it might seem surprising that these questions are not easily addressed today. This void is often filled by competing corporate marketing claims. For example, claims about risk detection model accuracy were discussed at the International Association of Gaming Regulators annual conference (percy2024trust), with examples including claims that models can exceed 80% accuracy (sustainableinteraction), 87% (mindwayai), and over 90% (entain2022sustain).
@@
Reviewing published academic papers or patent filings provides additional data beyond marketing and website claims but does not provide the full detail necessary (MARIONNEAU2025100644). While additional metrics and methodological detail are available, it is rare to find exactly the same reporting criteria across studies, simply because there are so many metrics and methodological choices involved. More seriously, even where full detail is available across contexts, there are no methods to adjust resulting performance metrics for true like-for-like comparison across those different contexts. Generalizability in player risk profiles and model structures is limited across brands (ghaharian2024evaluating), across time (Murch2024time), and likely across a host of other factors such as gambling type, jurisdiction and regulatory context. As a result, it is insufficient to lean solely on better norms for consistent reporting [^2] For examples of standardized reporting guidelines and checklists for ML models and AI-based systems, see TRIPOD-AI (Collins2024), CONSORT-AI (Liu2020), SPIRIT-AI (CruzRivera2020).. More transparent and standardized reporting is only one step towards answering a reasonable, fundamental question from end users: How does one model’s performance truly compare to others?
@@ 2.3 System Consequences
To be clear, these existing benchmarks are not perfect and are themselves the subject of ongoing research to improve their design (bean2025measuring; Hardy2024). But they prove that useful, non-simplistic comparisons are possible. Such comparisons support sector decisions and may fuel model improvements across the board, as they have done in some AI sectors, while building confidence and trust with stakeholders. At present no such benchmarks exist for player risk detection systems, so the only way of knowing which risk model works best in a given context is to build, buy, and apply them all—a profoundly impractical option.
@@ 3 Benchmark Datasets and Performance Benchmarking
A fundamental practice in ML research is evaluating algorithms against standardized benchmark datasets (longjohn2024benchmark; Thiyagalingam2022). Across numerous domains, benchmark datasets have proven successful in identifying optimal ML solutions for specific problems by enabling cross-stakeholder collaboration and fostering competitive environments for innovation. This process, Performance Benchmarking, allows for objective quantitative comparisons between different approaches, as methods and results cannot be reliably compared when models are trained on different data and evaluated with different diagnostics (Thiyagalingam2022). Such standardization helps track progress over time, assess novel methods, and select the most appropriate approach for a given application. Importantly, benchmark datasets should serve as proxies for real-world tasks, so performance on the benchmark is generalizable to practical applications (longjohn2024benchmark). Beyond scientific validation, benchmarking can also serve as a critical component of regulatory infrastructure, providing a transparent and accountable framework for evaluating complex systems (Percy2022). To understand how this framework could be applied in the gambling field, it is instructive to review its successful implementation in other domains.
@@ 3.1 Background
One of the earliest and most well-known examples of a benchmark dataset is the Modified National Institute of Standards and Technology (MNIST) database (lecun2010convolutional). Created in 1998, MNIST consists of 70,000 size-normalized, 28x28 pixel images of handwritten digits. Each image is accompanied by a label indicating the correct number. Before MNIST, tracking progress on numeric recognition was difficult because researchers used different and incompatible datasets. MNIST provided a large, standardized collection that allowed researchers to test and benchmark their models against a consistent baseline.
Throughout the 2000s, numerous benchmark datasets emerged, applying MNIST’s fundamental principles to a growing range of use cases and more complex applications. For example, the Street View House Number (SVHN) dataset, created in 2011, contains 600,000 labeled digits cropped from Street View images (netzer2011reading), and sought to advance image recognition beyond simple handwritten digits by providing a platform for identifying characters in natural environments. Similarly, the MedMNIST benchmark was created to advance biomedical image classification in healthcare (Yang2023).
More recently, the rapid advancements of LLMs have given rise to numerous benchmarks to assess their capabilities. Many, like the Massive Multitask Language Understanding (MMLU) test [^5] MMLU is designed to measure knowledge across diverse domains (e.g., humanities, social sciences, hard sciences) by evaluating a model’s accuracy on multiple-choice questions; for details, see its dataset card at https://huggingface.co/datasets/cais/mmlu., consist of large sets of multiple-choice questions across various domains (https://doi.org/10.48550/arxiv.2009.03300). Similar to how SVHN advanced image recognition, newer LLM benchmarks move beyond multiple-choice questions to assess more practical skills. For example, Vending-Bench uses a simulated environment to evaluate LLM-agents’ ability to operate a vending machine business (backlund2025vending) [^6] Vending-Bench is a simulated environment that tests how well an AI agent can manage a long-running business scenario. The agent must make decisions about inventory, orders, and pricing in order to make money (https://andonlabs.com/evals/vending-bench)..

### v2

@@ 1 Introduction
Gambling is a significant and growing form of entertainment, participated in by millions globally. While most consumers enjoy gambling as a leisure activity, prevalence rates for harm vary considerably across different populations, jurisdictions, and product types. Globally, an estimated 1.4% of adults meet criteria for problem gambling, but these rates shift dramatically depending on the gambling environment; for example, prevalence rises to approximately 16% among online casino or slots gamblers (Tran et al., 2024). The industry’s rapid expansion, driven largely by digitization and the global availability of online and mobile products, has led to a corresponding increase in efforts to mitigate these potential harms (Ghaharian et al., 2022). These collective efforts are commonly referred to as Responsible Gambling (RG) [^1] We use the term Responsible Gambling (RG) throughout this paper as it is commonly used to describe these practices. We acknowledge the ongoing and important discourse contrasting the RG framework to safer gambling or public health approaches, but a full discussion of that debate is beyond the scope of this paper..
@@
RG strategies have evolved as a direct result of gambling’s digitization, and include targeted RG messaging, immediate access to gambling help resources, and automated risk detection systems (Shaffer et al., 2020). These technological player protection measures can be divided roughly into two types: user-initiated measures and automated operator-initiated strategies. The former includes tools that users can employ, usually voluntarily, to control their gambling: self-limiting features, reality checks, deposit limits, cool-off periods, and self-exclusion. However, few users take advantage of these features (Ladouceur et al., 2016; Nelson et al., 2008), and there is debate about whether those who do are the players who need them most (Catania and Griffiths, 2021).
@@
While automated, operator-initiated strategies can also take many forms, including universal approaches like deposit limits, there has been a movement in the past decade toward more tailored interventions (Heirene and Wohl, 2025). Universal strategies risk either disrupting the activity of those who do not need protections or, if implemented in a way that does not disrupt most users’ activity, being set so conservatively that many of those who need them do not trigger them (Broda et al., 2008). On the other hand, targeted strategies can, in theory, use player information to better determine when and to whom to deploy protections.
@@
As such, data-driven RG models have become the holy grail of automated operator-initiated RG solutions (Ghaharian et al., 2022). These systems often leverage artificial intelligence (AI) – more specifically, machine learning (ML) – to analyze players’ gambling behavior and detect patterns associated with harm to trigger interventions. This approach has attracted substantial interest from both academia and industry stakeholders, with several regulators mandating the use of data-driven systems for player risk detection, operators developing proprietary models, and specialist technology vendors emerging to offer commercial solutions.
@@
The paper is presented as follows: We first provide a brief background on the use of ML for player risk detection, followed by a discussion of the current challenges in evaluating these systems. We then introduce the concept of Performance Benchmarking and briefly review its application in other domains. Finally, we present a conceptual, discussion-starting framework for a player risk detection benchmarking suite. Rather than proposing a single, fully specified benchmark, we outline the structural requirements, core dimensions, and practical considerations necessary for such benchmarks to be developed and adopted by the research and practitioner communities. Our hope is to stimulate discussion and thought for what such a framework would require and what it could achieve.
@@ 1.1 Machine Learning for Player Risk Detection
A ML approach to player risk detection typically uses information about gambling patterns and player characteristics from gambling records (behavioral markers) to forecast potential gambling-related problems [^2] For an up-to-date and comprehensive review of model architectures and the current landscape, we refer the reader to Marionneau et al., 2025. This is similar to the use of biomarkers in medicine, where objectively measured characteristics serve as indicators of disease processes (Biomarkers Definitions Working Group, 2001). Medicine also uses behavioral markers, for example, in early detection of Alzheimer’s disease (Gomar et al., 2011).
@@
ML models can target players according to markers defined by both their gambling patterns and changes in those patterns, using this information to designate differing levels of potential risk. Most importantly, these models can provide a guide for matching responses to individuals, optimizing the opportunity for favorable outcomes. Very basic best practices for developing such models include establishing that the markers differ in the groups to be distinguished (e.g., individuals who exhibit problematic gambling behavior and those who gamble without problems), developing a predictive model that integrates markers to predict the outcome in question, and then establishing the model’s performance in samples separate from those used to develop the model (Morrow et al., 2008; Vanderstichele et al., 2008).
@@
One distinct feature of ML is its ability to learn complex, nonlinear relationships in data, making it highly suited for identifying at-risk play given the individual variability in gambling behavior (Hastie et al., 2009). However, complex ML solutions encounter some of the same challenges as the most basic models, such as establishing accurate outcome measures and training data and developing models that can generalize and replicate beyond the idiosyncrasies of a specific dataset (i.e., distinguishing the signal from the noise: Silver, 2012; also see Shaffer et al., 2020).
@@
Early player risk models re-purposed data from loyalty smart cards used in physical gambling venues (e.g., casinos) (Schellinck and Schrans, 2011). The shift to online gambling, however, significantly simplified behavioral tracking. Unlike physical venues that depended on a player inserting a card, online platforms and mobile apps inherently capture comprehensive player data. This seamless data collection fueled a new research area, leading to at least 40 academic studies and several literature reviews (we refer the interested reader to: Ghaharian et al., 2022; Delfabbro et al., 2023; Marionneau et al., 2025). Alongside this academic work, commercial models were developed. A recent review by Marionneau et al. (2025) identified 11 commercial models, including six in-house systems from operators (e.g., Kindred, Veikkaus) and five from third-party vendors (e.g., Mindway, Openbet).
@@
These models vary widely in their methods, employing a variety of unsupervised and supervised algorithms, and in the number and type of input markers they use. As Marionneau et al. (2025) note, these markers are often unclear or inconsistently reported in the literature, and the number of markers used in a given model range from fewer than 10 to over 100. Ghaharian et al. (2025b) recently mapped the markers in the academic literature, identifying 65 “low-level” categories (such as bet amount and net loss) which can be grouped into five “high-level” categories: play (betting behavior), payment (financial transactions, i.e., deposits and withdrawals), engagement (when, how often, and how broadly players engage with gambling), RG tool use, and profile information (e.g., registration date, demographic information).
@@ 2 Current Problems
It is appropriate to hold ML to high standards, given that policy increasingly references it in supporting harm reduction (Bedford and Chandan, 2024). Moreover, such standards align with the sensible and realistic performance beliefs that gamblers themselves may hold (Murch et al., 2025). Several challenges limit the potential for ML in this context.
@@ 2.1 Broad Challenges for Stakeholders
ML, as with all identification and intervention design, involves trade-offs between sensitivity and precision, i.e., where to place the balance between failing to identify gamblers at risk and falsely identifying many gamblers with negligible risk levels (Murch et al., 2023). These trade-offs can be improved with more sophisticated models but cannot be fully resolved in real-world settings. In the absence of precision, an ultra-low specificity model simply provides RG interventions to all players, which may undermine trust in RG programs and potentially lead to substitution by players towards other operators or products.
@@
As models grow more complex, they typically become less transparent with respect to how any individual decision has been reached (Sarkar et al., 2016). This is often referred to as the “black box” problem, where the model’s internal logic is so intricate that it becomes opaque even to its creators, making it difficult to audit or explain a specific outcome (von Eschenbach, 2021). Explainability, or the ability to identify how a model arrives at its prediction, is important to facilitate users’ trust in a system and ensure the model is operating as expected (Sarkar et al., 2016). In addition, explainability has practical relevance: it enables operators to activate appropriate interventions and allows players to reflect on behavioral dimensions that have been identified as increasing their risk.
@@
Another issue is that models can only identify patterns in the data they possess. Consequently, identical gambling behaviors can represent very different risk-levels depending on variables not included in a data set, such as an individual player’s life circumstances or financial context. For example, the potential impact of a specific monetary loss is difficult to contextualize without information about an individual’s financial circumstances (e.g., income, debt exposure, etc.). Moreover, harms are not purely financial. A significant amount of time spent gambling could be a harmless hobby for one person, but for someone suffering psychological distress, it could be a coping mechanism leading to other harms like relationship disruption or missing work. Much of this relevant contextual information simply cannot be collected by operators under current social norms. This missing context is why seemingly identical profiles based on available data could have different risks and, as the UK Gambling Commission (2023) has noted, why high-risk players exist at all stake levels.
@@
There are also challenges related to AI ethics generally, such as risks of bias in datasets or models (Percy et al., 2020), data collection protocols, and ensuring human oversight for high-consequence decisions. While a full discussion of AI ethics in gambling is beyond this paper’s scope (discussions available in Binesh and Ghaharian, 2025; Ghaharian et al., 2025a; Percy et al., 2022), we note the trend towards translating ethical principles into regulation (e.g., the EU AI Act approved in May 2024).
@@ 2.2 Specific Challenges for Procuring or Commissioning Models
Given the quantitative nature of data science, it might seem surprising that these questions are not easily addressed today. This void is often filled by competing corporate marketing claims. For example, claims about risk detection model accuracy were discussed at the International Association of Gaming Regulators annual conference (Percy, 2024), with examples including claims that models can exceed 80% accuracy (Sustainable Interaction, n.d.), 87% (Mindway AI, n.d.), and over 90% (Entain, 2022).
@@
Reviewing published academic papers or patent filings provides additional data beyond marketing and website claims but does not provide the full detail necessary (Marionneau et al., 2025). While additional metrics and methodological detail are available, it is rare to find exactly the same reporting criteria across studies, simply because there are so many metrics and methodological choices involved. More seriously, even where full detail is available across contexts, there are no methods to adjust resulting performance metrics for true like-for-like comparison across those different contexts. Generalizability in player risk profiles and model structures is limited across brands (Ghaharian et al., 2024), across time (Murch et al., 2024b), and likely across a host of other factors such as gambling type, jurisdiction and regulatory context. As a result, it is insufficient to lean solely on better norms for consistent reporting [^3] For examples of standardized reporting guidelines and checklists for ML models and AI-based systems, see TRIPOD-AI (Collins et al., 2024), CONSORT-AI (Liu et al., 2020), SPIRIT-AI (Cruz Rivera et al., 2020).. More transparent and standardized reporting is only one step towards answering a reasonable, fundamental question from end users: How does one model’s performance truly compare to others?
@@ 2.3 System Consequences
To be clear, these existing benchmarks are not perfect and are themselves the subject of ongoing research to improve their design (Bean et al., 2025; Hardy et al., 2024). But they prove that useful, non-simplistic comparisons are possible. Such comparisons support sector decisions and may fuel model improvements across the board, as they have done in some AI sectors, while building confidence and trust with stakeholders. At present no such benchmarks exist for player risk detection systems, so the only way of knowing which risk model works best in a given context is to build, buy, and apply them all—a profoundly impractical option.
@@ 3 Benchmark Datasets and Performance Benchmarking
A fundamental practice in ML research is evaluating algorithms against standardized benchmark datasets (Longjohn et al., 2024; Thiyagalingam et al., 2022). Across numerous domains, benchmark datasets have proven successful in identifying optimal ML solutions for specific problems by enabling cross-stakeholder collaboration and fostering competitive environments for innovation. This process, Performance Benchmarking, allows for objective quantitative comparisons between different approaches, as methods and results cannot be reliably compared when models are trained on different data and evaluated with different diagnostics (Thiyagalingam et al., 2022). Such standardization helps track progress over time, assess novel methods, and select the most appropriate approach for a given application. Importantly, benchmark datasets should serve as proxies for real-world tasks, so performance on the benchmark is generalizable to practical applications (Longjohn et al., 2024). Beyond scientific validation, benchmarking can also serve as a critical component of regulatory infrastructure, providing a transparent and accountable framework for evaluating complex systems (Percy et al., 2022). To understand how this framework could be applied in the gambling field, it is instructive to review its successful implementation in other domains.
@@ 3.1 Background
One of the earliest and most well-known examples of a benchmark dataset is the Modified National Institute of Standards and Technology (MNIST) database (LeCun et al., 2010). Created in 1998, MNIST consists of 70,000 size-normalized, 28x28 pixel images of handwritten digits. Each image is accompanied by a label indicating the correct number. Before MNIST, tracking progress on numeric recognition was difficult because researchers used different and incompatible datasets. MNIST provided a large, standardized collection that allowed researchers to test and benchmark their models against a consistent baseline.
Throughout the 2000s, numerous benchmark datasets emerged, applying MNIST’s fundamental principles to a growing range of use cases and more complex applications. For example, the Street View House Number (SVHN) dataset, created in 2011, contains 600,000 labeled digits cropped from Street View images (Netzer et al., 2011), and sought to advance image recognition beyond simple handwritten digits by providing a platform for identifying characters in natural environments. Similarly, the MedMNIST benchmark was created to advance biomedical image classification in healthcare (Yang et al., 2023).
More recently, the rapid advancements of LLMs have given rise to numerous benchmarks to assess their capabilities. Many, like the Massive Multitask Language Understanding (MMLU) test [^6] MMLU is designed to measure knowledge across diverse domains (e.g., humanities, social sciences, hard sciences) by evaluating a model’s accuracy on multiple-choice questions; for details, see its dataset card at https://huggingface.co/datasets/cais/mmlu., consist of large sets of multiple-choice questions across various domains (Hendrycks et al., 2020). Similar to how SVHN advanced image recognition, newer LLM benchmarks move beyond multiple-choice questions to assess more practical skills. For example, Vending-Bench uses a simulated environment to evaluate LLM-agents’ ability to operate a vending machine business (Backlund and Petersson, 2025) [^7] Vending-Bench is a simulated environment that tests how well an AI agent can manage a long-running business scenario. The agent must make decisions about inventory, orders, and pricing in order to make money (https://andonlabs.com/evals/vending-bench)..